OpenStudy (anonymous):

what is the simplified form of x squared plus 5 divided by x plus 10

OpenStudy (ganpat):

(x2+5)/(x+10) = 1 x2 + 5 = x + 10 can u solve further ?

OpenStudy (ganpat):

bring all the variables on one side...

OpenStudy (anonymous):

No i cant can u solve for me please so i can compare to my answer

OpenStudy (ganpat):

whats your answer ?

OpenStudy (ganpat):

x2 -x +5 - 10 = 0 x2 - x - 5 = 0